OpenStudy (anonymous):

What is the arc length if Θ = 6 pi over 5 and the radius is 2 cm? 3 pi over 10 24 pi over 5 5 pi over 12 12 pi over 5

OpenStudy (anonymous):

12pi/5 ithink?

OpenStudy (anonymous):

s=theta*radius s=arc length

OpenStudy (anonymous):

true 12pi over 5 cause the formula is ark length =angle x radius

OpenStudy (anonymous):

If sin Θ = 12 over 37, use the Pythagorean Identity to find cos Θ. cos Θ = ± 35 over 37 cos Θ = ± 23 over 37 cos Θ = ± 35 over 12 cos Θ = ± 20 over 35

OpenStudy (anonymous):

35/37

OpenStudy (anonymous):

sin theta= opp/hyp thus cos= adj/hyp using pythagorean theorem we know x^2+12^2=37^2 solving for x we have=> x^2+144=1369 x^2=1225 x=sqrt(1225) =>37 answer: 35/37